OpenStudy (anonymous):

-45r^4+42r^3-12r^2+14r+5/-3r Simplify

OpenStudy (mertsj):

\[\frac{-45r^4+42r^3-12r^2+14r+5}{-3r}=15r^3-14r^2+4r+\frac{14r+5}{-3r}\]